a) What is the energy (in J) stored in the 23.0 µF capacitor of a heart defibrillator charged to 8.90 ✕ 103 V? J (b) Find the amount of stored charge (in mC). mC
Answer #1

a) E = 1/2 C * V^2
E = 0.5*23*10^-6*(8.9*10^3)^2 = 910.915 Joule

b ) Q = C V = 23*10^-6 * 8900 = 0.02047 Coulomb = 20.47 mC